Part 1.C&A Fast Food has four activities in serving a customer: greet customer, take order, process order, and deliver order. Each activity is staffed by one employee (for a total of four employees). The processing time for each activity is given as follows:

ActivityProcessing time per customer

Greet customer 2 seconds

Take order 30 seconds

Process order 60 seconds

Deliver order 5 seconds.

a. What is the average labor utilization?

b.Assume demand is unlimited. If one additional employee is added to the bottleneck activity, what will be the percentage change in process capacity?

Part 2.Which of the following statements about takt time and cycle time is TRUE ?

A) If takt time is longer than cycle time, target manpower will be lower than current staff level.

B) Supply matches demand when takt time is shorter than cycle time.

C) If takt time is longer than cycle time, target manpower will be higher than current staff level.

D) Supply is less than demand when takt time is equal to cycle time.
Answer #1

Part 1

a.In this case the bottleneck is the step “Process order” with a capacity of 1 customer per minute. This will reduce the overall capacity of the system. At this rate the system capacity is 1 customer per minute. This also results in the usage of the following times = 2 + 30 + 60 + 5 = 97 seconds. However, the total available time is 4*60 = 240 seconds. This means the utilization is 97/240 = 0.4041 or 40.41%

b.If we add another employee to “Process order” the capacity of the step will double. This means step capacity will be 2 customer per minute. This is also the same as “Take order” step. Overall process capacity will now be 2 customers per minute. This is a 100% increase over the previous stage.

Part 2

If takt time is longer than cycle time, target manpower will be lower than current staff level.

Takt time represents the time the company should take to make a product. Cycle time represents the time company takes to make a product. This means if the cycle time is less or takt time is high, the company can meet the demand and also have some idle time. This translates into the fact that the current manpower is more than target manpower.